OpenStudy (anonymous):

is the power greater when 500 J of work is done in 5 seconds or in 10 seconds

OpenStudy (anonymous):

I WILL FAN AND MEDAL

OpenStudy (anonymous):

\[P =\frac{ \Delta W }{ \Delta t }\]

OpenStudy (shamim):

P1=500/5=? P2=500/10=?

OpenStudy (shamim):

Which is greater p1or p2

OpenStudy (anonymous):

oh its p1